# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

### v1

@@ III-B FL Framework Mechanism Adopted
Here, we consider a set of =$\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs with the capacities of local computation and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^6] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive updated models from all MUs before performing model aggregation. The advantages of synchronous FL are high model accuracy and fast convergence [22, 23].. Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}(n\in\mathcal{N})$. The MUs can join FL to generate local learning models and contribute to enhancing vehicular metaverse services. As shown in Fig. 4, MU $m$ provides learning models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel. Moreover, the data collected at the $(r-1)$-th round is used as the local data for training at the $r$-th round. Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of the period $T$.
Figure 4: An illustration of the FL mechanism with $n$ tasks for MU $m$.
To be concrete, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$. The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles. The steps involved in each iteration $r\in\left\{1,2,\ldots,R\right\}$ are as follows:
@@
    $\jmath_{n}^{(r+1)}$ $=\jmath^{(r)}_{n}+\sum_{m=1}^{M}\frac{|X_{mn}^{(r)}|}{|X_{n}^{(r)}|}(\jmath_{n}^{(r)}-\jmath_{mn}^{(r)}),$ (1)
@@ III-C Immersion Metric of Learning Model
MSPs determine the rewards for MUs based on the immersion of the learning model (IoM) metric. Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computation and communication resources to maximize the benefits. This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making. Unlike traditional FL and incentive mechanisms, in vehicular metaverse scenarios with immersion requirements, both the contribution and freshness of the learning model need to be considered [^7] For example, an MSP may receive fresher learning models but with lower contributions, or less fresh learning models but with higher contributions. , which will collectively affect the immersive experience of AR services, i.e., whether virtual objects can be placed in the physical world accurately and promptly. With the above consideration, we propose the IoM, a metric that jointly considers the contribution prediction $I_{mn}$ and the age of information (AoI) $\Delta_{mn}$ [24] to measure the value of a learning model that MU $m$ brought to the vehicular metaverse service of MSP $n$, denoted as
@@ III-C1 Contribution Prediction $I$
$I_{mn}=\frac{\omega_{mn}\epsilon ln(1+\eta\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|)}{\theta_{m}},$ (3)
@@ III-C2 Age of Information $\Delta$
which shows the elapsed time since the generation of the latest updated model. A smaller instantaneous AoI means that the latest model received is fresher at that time. To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models. With the aid of Figure 5, the average AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
@@
Since MU $m$ constantly collects data throughout the FL process, the amount of training data collected for training in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time. The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computation resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
@@
where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-interference-plus-noise ratio (SINR) for the communication between MU $m$ and MSP $n$. $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively. $\sigma^{2}$ is additive white Gaussian noise (AWGN).
@@ IV-A Utility of MU
We define the utility of MU $m$ composed of the rewards from MSPs and the cost incurred for contribution, which is also affected by resource constraints and immersion requirements for consuming vehicular metaverse services. The cost incurred due to local training and model uploading, which can be expressed as
@@
where $c_{m}^{f}$ and $c_{m}^{B}$ denote the cost factors of computation and communication resources, respectively. Let $\bm{f}_{m}\triangleq\left(f_{m1},f_{m2},\ldots,f_{mn}\right)$ and $\bm{B}_{m}\triangleq\left(B_{m1},B_{m2},\ldots,B_{mn}\right)$ be the computation and communication allocation profiles of MU $m$, respectively. Then, the utility maximization problem for MU $m$ within $T$ can be formulated as
Problem 1
@@
where $S_{m}$ indicates the minimum computation resource required to enjoy other basic services. Constraints C1 and C3 indicate that the MU’s total computation and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services. Constraint C2 ensures that the total time for local training and uploading cannot exceed the time constraint $\tau_{n}$ for each round. The goal of each MU is to choose the optimal allocation of computation $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility. Note that $V_{mn}$ must be greater than 00. This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, then the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ will not participate in the trading.
@@ IV-B Utility of MSP
Each MSP has a gain function $\psi$ associated with the IoM to measure the benefits of all the learning models from MUs to the service. In this paper, we adopt $\psi_{n}=\mu_{n}ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [28]. Let $\bm{p}_{n}\triangleq\left(p_{1n},\ldots,p_{mn}\right)$ and $\bm{p}_{\textrm{-}n}$ represent the reward profile of MSP $n$ and all other MSPs except MSP $n$, respectively. Then, the optimization problem for the MSP is defined as
Problem 2
@@
where $\mu_{n}$ refers to the profit conversion coefficient [29] from IoM, which is adjusted according to the AR services offered by the different MSPs. Note that the other MSPs’ decisions $\bm{p}_{\textrm{-}n}$ are captured by the MUs’ responses. Moreover, when MSP $n$ determine its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$). This thereby leads to reward competition among MSPs. Therefore, the optimization among MSPs can be considered as a non-cooperative game, termed a multi-MSP rewarding game $\Omega$ as follows.
@@
Figure 6: The structure of EPEC.
@@ IV-C Multi-MSP Multi-MU Game as EPEC
In such a multi-MSP and multi-MU scenario, each MU has the choice of multiple MSPs to contribute learning models, and each MSP can incentivize multiple MUs to earn high-value learning models and achieve more benefits. Accordingly, each MSP competes with others for the equilibrium, which is constrained by the equilibrium among the MUs. This leads to the EPEC formulation [30].
As shown in Fig. 6, MUs and MSPs negotiate allocation strategies of computation and communication resources and reward decisions to maximize their benefits. In the upper level, MSPs determine the rewards they are willing to offer by considering their cost, the responses of MUs, and the decisions of other MSPs. In the lower level, each MU m gives the optimal allocation response for computation and communication resources by considering its resource constraints, cost, and rewards from different MSPs. The objective of EPEC is to find the equilibria at two levels, i.e., the point at which the MSPs’ (leaders’) utilities are maximized given that the MUs (followers) will choose their best responses. For the proposed EPEC, the equilibria at two levels are defined as follows.
@@
Let $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ denote the optimal computation and communication resource allocation of MU $m\in\mathcal{M}$ and the optimal reward decision of MSP $n\in\mathcal{N}$, respectively. Then, the points $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ are the equilibria at two levels if the following conditions hold:
@@
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N}.$
@@
V Analysis and Solutions
In this section, we use backward induction to investigate the existence and uniqueness of the equilibria at two levels, i.e., the optimal computation and communication resource allocation of MUs at the lower level and the optimal reward decisions of MSPs at the upper level.
V-A Lower Level: Optimal Computation and Communication Resource Allocation Analysis
For any reward decisions $\bm{p}$ given by MSPs, MU $m$ would like to determine an optimal decision $\bm{f}_{m}^{*}$ and $\bm{B}_{m}^{*}$ to maximize its utility. We analyze and provide the solution for the optimal allocation of computation and communication resources as follows.
@@ V-A Lower Level: Optimal Computation and Communication Resource Allocation Analysis
If the problem is strictly concave for maximization and satisfies the following conditions, it will have a unique optimal solution [31]: (1)The Utility function is concave; (2) Constraints are all affine constraints; and (3) Strategy space is a compact and convex set.
@@
Problem 1 has a unique optimal solution in the lower level, i.e., each MU has an optimal and unique allocation scheme for computation and communication resources.
@@
We first observe that the Hessian matrix of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $(f_{mn},B_{mn})$ is
$\bm{H}_{m}=\begin{bmatrix}\bm{H}_{m}^{f}&0\\ 0&\bm{H}_{m}^{B}\end{bmatrix},$ (15)
where $\bm{H}^{f}_{m}$ can be obtained by the second-order derivation of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $f_{mn}$
@@
where $h^{f}_{mn}=p_{mn}I_{mn}\frac{2x_{m}\tau_{n}}{f_{mn}^{3}}\log(1/\theta_{m})$.
Similarly, the $\bm{H}^{B}_{m}$ can be calculated as
@@
We can easily show that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are negative semi-definite, and thus $\bm{H}_{m}$ is negative semi-definite and $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is concave. Moreover, according to the convexity condition of its constraints and Definition 3, it can be proved that Problem 1 has a unique optimal solution $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$. ∎
Next, we can obtain the best response function of MU $m$ by solving Problem 1. Based on the first-order condition and setting the derivative to zero, we can derive the following proposition.
Proposition 1. Given the digital currency price offered by the MSPs and the Lagrange multipliers induced by the constraints, the best response function $f_{mn}$ of MU $m$ satisfies:
@@
and the best response function $B_{mn}$ of MU $m$ satisfies:
$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
where $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}$.
Proof.
First, we transform Problem 1 into an unconstrained optimization problem using the Lagrangian dual method. The Lagrangian function associated with MU $m$ is formulated as
$-\sum_{n}\left[c_{m}^{f}f_{mn}log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
$-\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})-\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})$
$-\sum_{n}\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]$
$-\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})],$
$\frac{\partial L_{m}}{\partial f_{mn}}$ $=p_{mn}I_{mn}\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}-c_{m}^{f}log(1/\theta_{m})$ (21)
$-\lambda_{m}+\delta_{mn}log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-\gamma_{m}T^{req},$
and
$\frac{\partial L_{m}}{\partial B_{mn}}$ $=p_{mn}I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}-\beta_{m}$ (22)
$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}.$
From (21) and the constraints in (11), the KKT conditions are given by
$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
$\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})]=0,$
$\lambda_{m},\delta_{mn},\gamma_{m}\geq 0,C1,C2,C3.$
$(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-c_{m}^{f}=0.$ (24)
Similarly, the KKT conditions for $B_{mn}$ are obtained as
$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
$\beta_{m},\delta_{mn}\geq 0,C1,C2.$
$(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}=0.$ (26)
  • ($\delta_{mn}=0$): According to (24) and (26), we have
    $f_{mn}=\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (27)
  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}\right]=0$, we can get
    $\delta_{mn}=F^{2}-p_{mn}I_{mn},$ (28)
    $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}.$
    Next, bring $\delta_{mn}$ into $f_{mn}$ and $B_{mn}$, we have
    $f_{mn}=F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (29)
According to Proposition 1, the optimal strategy for each MU $m$ is influenced by five main factors: $p_{mn}$, $I_{mn}$, $c_{m}^{f}$, $c_{m}^{B}$ and $\tau_{n}$. Specifically, MU $m$ focuses on the rewards paid by MSPs, and the MU tends to put more computation and communication resources into the contribution when $p$ increases. Moreover, MUs are more willing to invest more resources when the potential contribution prediction $I_{mn}$ is higher, resulting in greater benefits. In addition, the virtual deadline $\tau_{n}$ set by the MSP $n$ determines the minimum criteria for resource allocation.
V-B Upper Level: Optimal Reward Decision Equilibrium Analysis
Given the responses $(\bm{f},\bm{B})$ from all MUs, and other MSPs’ decisions $\bm{p}_{\textrm{-}n}$, the optimal reward decision $\bm{p}_{n}$ of MSP $n$ can be obtained by solving Problem 2, defined as follows.
Proposition 2. Given $\bm{p}_{\textrm{-}n}$, MSP $n$’s best response strategy is

### v2

@@ III-B FL Framework Mechanism Adopted
Here, we consider a set of =$\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs with the capacities of local computation and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^7] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive updated models from all MUs before performing model aggregation. The advantages of synchronous FL are high model accuracy and fast convergence [29].. Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}(n\in\mathcal{N})$. The MUs can join FL to generate local learning models and contribute to enhancing vehicular metaverse services. As shown in Fig. 4, MU $m$ provides learning models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel. Moreover, the data collected at the $(r-1)$-th round is used as the local data for training at the $r$-th round. Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of the period $T$.
Figure 4: Illustration of the FL mechanism with $n$ tasks for MU $m$.
In detail, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$. The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles. The steps involved in each iteration $r\in\left\{1,2,\ldots,R\right\}$ are as follows:
@@
    $\jmath_{n}^{(r+1)}$ $=\jmath^{(r)}_{n}+\sum_{m=1}^{M}\frac{|X_{mn}^{(r)}|}{|X_{n}^{(r)}|}(\jmath_{mn}^{(r)}-\jmath_{n}^{(r)}),$ (1)
@@ III-C Immersion Metric of Learning Model
MSPs determine the rewards for MUs based on the immersion of the learning model (IoM) metric. Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computational and communication resources to maximize the benefits. This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making. Unlike traditional FL and incentive mechanisms, in vehicular metaverse scenarios with immersion requirements, both the contribution and freshness of the learning model need to be considered [^8] For example, an MSP may receive fresher learning models but with lower contributions, or less fresh learning models but with higher contributions. , which will collectively affect the immersive experience of AR services, i.e., whether virtual objects can be placed in the physical world accurately and promptly. With the above consideration, we propose the IoM, a metric that jointly considers the contribution prediction $I_{mn}$ and the age of information (AoI) $\Delta_{mn}$ [30] to measure the value of a learning model that MU $m$ brought to the vehicular metaverse service of MSP $n$, denoted as
@@ III-C1 Contribution Prediction $I$
$I_{mn}=\frac{\omega_{mn}\epsilon\ln(1+\eta\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|)}{\theta_{m}},$ (3)
@@ III-C2 Age of Information $\Delta$
which shows the elapsed time since the generation of the latest updated model. A smaller instantaneous AoI means that the latest model received is fresher at that time. To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models. With the aid of Figure 5, the time-averaged AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
@@
Since MU $m$ constantly collects data throughout the FL process, the amount of training data collected for training in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time. The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computational resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
@@
where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-noise ratio (SNR) [34] for the communication between MU $m$ and MSP $n$. $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively. $\sigma^{2}$ is additive white Gaussian noise (AWGN).
@@ IV-A Utility of MU
We define the utility of MU $m$ composed of the rewards from MSPs and the cost incurred for contribution, which is also affected by resource constraints and immersion requirements for consuming vehicular metaverse services. The cost incurred due to local training and model uploading can be expressed as
@@
where $c_{m}^{f}$ and $c_{m}^{B}$ denote the cost factors of computational and communication resources, respectively. Let $\bm{f}_{m}\triangleq\left(f_{m1},f_{m2},\ldots,f_{mn}\right)$ and $\bm{B}_{m}\triangleq\left(B_{m1},B_{m2},\ldots,B_{mn}\right)$ be the computational and communication allocation profiles of MU $m$, respectively. Then, the utility maximization problem for MU $m$ within $T$ can be formulated as
Problem 1.
@@
where $S_{m}$ indicates the minimum computational resource required to enjoy other basic services. Constraints C1 and C3 indicate that the MU’s total computational and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services. Constraint C2 ensures that the total time for local training and uploading cannot exceed the time constraint $\tau_{n}$ for each round. The goal of each MU is to choose the optimal allocation of computational $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility. Note that $V_{mn}$ must be greater than 00. This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ choosing not to participate in the trading.
@@ IV-B Utility of MSP
Each MSP has a gain function $\psi$ associated with the IoM to measure the benefits of all the learning models from MUs to the service. In this paper, we adopt $\psi_{n}=\mu_{n}\ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [35]. Let $\bm{p}_{n}\triangleq\left(p_{1n},\ldots,p_{mn}\right)$ and $\bm{p}_{\textrm{-}n}$ represent the reward profile of MSP $n$ and all other MSPs except MSP $n$, respectively. Then, the optimization problem for the MSP is defined as
Problem 2.
@@
where $\mu_{n}$ refers to the profit conversion coefficient [36] from IoM, which is adjusted according to the AR services offered by the different MSPs. Note that the other MSPs’ decisions $\bm{p}_{\textrm{-}n}$ are captured by the MUs’ responses. Moreover, when MSP $n$ determines its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$). This thereby leads to reward competition among MSPs. Therefore, the optimization among MSPs can be considered as a non-cooperative game, termed a multi-MSP rewarding game $\Omega$ as follows.
@@
Figure 6: The hierarchical structure of equilibrium problem with equilibrium constraints (EPEC).
@@ IV-C Multi-MSP Multi-MU Game as EPEC
As shown in Fig. 6, MUs and MSPs negotiate allocation strategies of computational and communication resources and reward decisions to maximize their benefits. In the upper level, MSPs determine the rewards they are willing to offer by considering their cost, the responses of MUs, and the decisions of other MSPs. In the lower level, each MU $m$ gives the optimal allocation response for computational and communication resources by considering its resource constraints, cost, and rewards from different MSPs. The objective of EPEC is to find the equilibria at two levels, i.e., the point at which the MSPs’ (leaders’) utilities are maximized given that the MUs (followers) will choose their best responses. For the proposed EPEC, the equilibria at two levels are defined as follows.
@@
Let $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ denote the optimal computational and communication resource allocation of MU $m\in\mathcal{M}$ and the optimal reward decision of MSP $n\in\mathcal{N}$, respectively. Then, the points $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ are the equilibria at two levels if the following conditions hold:
@@
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N},$
where $\bm{p}_{-n}^{*}$ denotes the optimal rewards offered by other MSPs.
@@
V EPEC Analysis and Solutions
In this section, we utilize backward induction to analyze the EPEC formulated in (14). Specifically, we prove the existence and uniqueness of the equilibrium solutions at two hierarchical levels: (1) the optimal allocation of computational and communication resources for MUs (lower level), and (2) the optimal reward decisions for MSPs (upper level).
V-A Lower Level: Optimal Resource Allocation for MUs
In the lower level of EPEC, for any reward decisions $\bm{p}$ given by MSPs, MU $m$ aims to solve Problem 1 in (11) to determine its optimal computational and communication resource allocations, i.e., $\bm{f}_{m}^{*}$ and $\bm{B}_{m}^{*}$, to maximize its utility. Below, we analyze and derive the unique optimal allocation.
@@ V-A Lower Level: Optimal Resource Allocation for MUs
A maximization problem has a unique global optimum if the following conditions are satisfied [37]: (1) The objective function is strictly concave; (2) All constraints (equality and inequality) are affine; and (3) The strategy space is compact and convex.
@@
Problem 1 is strictly concave and has a unique globally optimal solution in the lower level. That is, for each MU $m$, there exists a unique resource allocation tuple $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$ that maximizes its utility.
@@
We first examine the Hessian matrix of MU $m$’s utility $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $(f_{mn},B_{mn})$. Let this Hessian matrix be denoted by $\bm{H}_{m}$, which can be block-diagonalized as
$\bm{H}_{m}=\begin{bmatrix}\bm{H}_{m}^{f}&0\\ 0&\bm{H}_{m}^{B}\end{bmatrix}.$ (15)
The block matrix $\bm{H}^{f}_{m}$ can be computed as the second-order partial derivative of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $f_{mn}$, i.e.,
@@
where $h^{f}_{mn}=p_{mn}I_{mn}\frac{2x_{m}\tau_{n}}{f_{mn}^{3}}\log(1/\theta_{m})$. Similarly, the block matrix $\bm{H}^{B}_{m}$ can be calculated by
@@
We find that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are both negative semi-definite. Hence, $\bm{H}_{m}$ is negative semi-definite, implying $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is strictly concave. Since the constraints of Problem 1 are affine and the strategy space is closed and bounded, Definition 3 ensures the existence of a unique global optimum $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$. ∎
Consequently, MU $m$ has a unique best-response resource allocation for each MSP by solving Problem 1. From the Karush–Kuhn–Tucker (KKT) conditions, we establish the following proposition.
Proposition 1. Given the digital currency price offered by the MSPs and the Lagrange multipliers induced by the constraints, the optimal computational resource $f_{mn}$ allocated for MSP $n$ by MU $m$ satisfies:
@@
and the optimal communication bandwidth $B_{mn}$ is:
$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
where $F=\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}$. For the detailed proof of Proposition 1, please refer to Appendix A.
According to Proposition 1, the optimal strategy for each MU $m$ is influenced by five main factors: $p_{mn}$, $I_{mn}$, $c_{m}^{f}$, $c_{m}^{B}$ and $\tau_{n}$. Specifically, MU $m$ focuses on the rewards paid by MSPs, and the MU tends to put more computational and communication resources into the contribution when $p$ increases. Moreover, MUs are more willing to invest more resources when the potential contribution prediction $I_{mn}$ is higher, resulting in greater benefits. In addition, the virtual deadline $\tau_{n}$ set by the MSP $n$ determines the minimum criteria for resource allocation.
V-B Upper Level: Optimal Reward Equilibrium among MSPs
In the upper level of the EPEC, each MSP $n$ competes with other MSPs and determines its reward vector $\bm{p}_{n}$. Specifically, given the responses $(\bm{f},\bm{B})$ from all MUs, and other MSPs’ decisions $\bm{p}_{\textrm{-}n}$, the optimal reward decision $\bm{p}_{n}$ of MSP $n$ can be obtained by solving Problem 2, defined as follows.
Proposition 2. Given other MSPs’ reward vectors $\bm{p}_{\textrm{-}n}$, the optimal strategy of MSP $n$ is